# High-Power Test of a C-band Linear Accelerating Structure with an RFSoC-based LLRF System

### v1

@@ IV.2 Rf pulse with phase reversal
Phase reversal applied to the rf pulse became common for pulse compressors since SLED has been implemented and delivered stable performance. From the first SLED implementation to more recent modified versions, the phase reversals were realized with a separate phase shifter or inverter components. For the NG-LLRF platform, the modulation processes are fully implemented in digital domain. A base-band waveform in any shape can be generated in software and modulated in the digital data-paths integrated in RFSoC. In this case, the rf signal has been modulated with a square wave with a phase reversal every 200 ns to understand the system behavior with phase reversal. The rf pulse delivered to the test structure in this case has the peak power of 4.2 MW and pulse duration approximately 1 $\mu$s at 10 Hz.
@@
In the first 200 ns after the rf power arrived the structure, the cavity has been filled in the same way as square wave. When the phase is reversed at 200 ns, the rf power has been extracted by the field out of phase rapidly. The extraction appears as the second peak followed by a rapid falling on the cavity reflection signal, which is over 160% higher than the first peak. This filling and extraction process cycle is similar to the method used by SLED system to produce a short pulse with higher power. After the field is completely extracted, the structure is filled with the rf field with reversed phase, which occurs as the slower decay after the rapid falling. When the phase is reversed again, the power extraction and filling repeats again, which corresponds to the third peak and delay followed. This cycle alternates until the rf power is down.
The phase reversal test has been performed with 450 ns rf pulses with peak power of 16.45 MW at 60 Hz. The rf signals follow similar trends as the signals measured with 4.2 MW and 1 $\mu$s rf pulses. As the power level is significantly higher, the cavity reflection signal declines with higher gradient and the field filling process is more obvious in the first 200 ns. In the power extraction and filling process after the second peak of cavity reflection signal, the change in the gradient shows the transition between the rapid field extraction and filling with inverted phase.
@@ IV.3 Rf pulse with linear phase sweep
The phase modulation technique has been used to reduced the rf power requirement and beam loading compensation of High-Luminosity Large Hadron Collider (HL-LHC) Mastoridis, Baudrenghien, and Molendijk (2017). The technique could be used for future particle accelerators for those purposes or beyond. In this case, the rf signal has been modulated with a linear phase ramp from -180°  to 180°  in 500 ns to understand the responses of the rf drive and the C3 structure to the phase ramp at peak power up to 16.45 MW at 60 Hz pulse rate.
Figure 14 shows the rf signals measured with rf pulse with 360°  linear phase ramp at peak power level of 16.45 MW. The three rf signals have different initial phase values, but the 360°  phase ramps are fully completed for all of them. The 360°  phase linear sweep in 500 ns can be also considered as driving the test setup with around 2 MHz frequency offset. The klystron has the bandwidth to generate the phase sweep. However, the accelerating structure cavity has bandwidth in kHz range and the 2 MHz offset is far from the resonance of the strucuture. As Figure 14 shows, the cavity forward and refection signal have extremely similar waveforms, which indicates that the rf power forwarded to the structure is almost fully reflected.
@@ V Conclusion
In this high-power test of the C3 accelerating structure driven by NG-LLRF, the amplitude and phase stability levels have been evaluated comprehensively at different stages of the test setup. The direct loop-back characterization test of the NG-LLRF prototype demonstrated around 70 fs phase jitters and 0.13 % of amplitude fluctuation. That is a critical milestone for the new direct rf sampling technique employed to implement the NG-LLRF platform. The C-band rf pulses are directly generated by the DACs and measured by ADCs without any analogue rf mixing in higher order Nyquist zones, which can significantly simplify the structure of a LLRF system. The NG-LLRF can be deployed rapidly for accelerators operated with rf frequency lower than 6 GHz.
@@
The high-power test with square rf pulse reveals the full filling and dissipation processes and other reflections between the rf components. On the cavity forward signal, there is a decline after the initial ramp up, which is the result of cross coupling between forward and reflection couplers. For accelerators operation with bunch trains, the requirement in the flatness of the pulse top is more stringent. Therefore, the couplers for measuring the forward signal should be designed to minimize the cross coupling and capture the forward signal with high accuracy. For the test of rf pulse with phase reversal, the high power short pulses have been successfully generated on the reflection signal. Therefore, the NG-LLRF can be employed to drive and control a SLED structure for pulse compression without any analogue modulator or phase shifter. For the tests of rf pulse with a linear phase sweep in 500 ns, the klystron generates the phase ramp with high precision, but rf power forwarded to the cavity has been fully reflected, as the phase ramp is equivalent of drive the structure off the resonance by approximately 2 MHz.

### v2

@@ IV.2 Rf pulse with phase reversal
Phase reversal applied to the rf pulse became common for pulse compressors since SLED has been implemented and it delivered stable performance. From the first SLED implementation to more recent modified versions, the phase reversals were realized with a separate phase shifter or inverter components. For the NG-LLRF platform, the modulation processes are fully implemented in digital domain. A base-band waveform in any shape can be generated in software and modulated in the digital data-paths integrated in RFSoC. In this case, the rf signal has been modulated with a square wave with a phase reversal every 200 ns to understand the behavior of the system with phase reversal. The rf pulse delivered to the test structure in this case has a peak power of 4.2 MW and a pulse duration of approximately 1 $\mu$s at 10 Hz.
@@
In the first 200 ns after the rf power arrived the structure, the cavity has been filled in the same way as it was driven by a square wave. When the phase is reversed at 200 ns, the rf power has been extracted by the field out of phase rapidly. The extraction appears as the second peak followed by a rapid fall on the cavity reflection signal, which is over 160% higher than the first peak. This filling and extraction process cycle is similar to the method used by the SLED system to produce a short pulse with higher power. After the field is completely extracted, the structure is filled with the rf field with reversed phase, which occurs as the slower decay after the rapid fall. When the phase is reversed again, the power extraction and filling repeat again, and that corresponds to the third peak and the decay followed. This cycle alternates until the rf power is down.
The phase reversal test has been performed with 450 ns rf pulses with peak power of 16.45 MW at 60 Hz. As Figure 13 shows, the rf signals captured at 16.45 MW follow trends similar to the signals measured at 4.2 MW shown in Figure 12. As the power level is significantly higher, the cavity reflection signal declines with higher gradient and the field filling process is more obvious in the first 200 ns. In the power extraction and filling process after the second peak of the cavity reflection signal, the change in the gradient shows the transition between rapid field extraction and filling with an inverted phase.
@@ IV.3 Rf pulse with linear phase sweep
The phase modulation technique has been used to reduce the rf power requirement and beam loading compensation of High-Luminosity Large Hadron Collider (HL-LHC) Mastoridis, Baudrenghien, and Molendijk (2017). The technique could be used for future particle accelerators for those purposes or beyond. In this case, the rf signal has been modulated with a linear phase ramp from -180°  to 180°  in 500 ns to understand the responses of the rf drive and the C3 structure to the phase ramp at peak power up to 16.45 MW at 60 Hz of pulse rate.
Figure 14 shows the rf signals measured with rf pulse with 360°  linear phase ramp at peak power level of 16.45 MW. The three rf signals have different initial phase values, but the 360°  phase ramps are fully completed for all of them. The 360°  phase linear sweep in 500 ns can also be considered as driving the test setup with around 2 MHz frequency offset. The klystron has the bandwidth to generate the phase sweep. However, the accelerating structure cavity has a bandwidth in kHz range and the 2 MHz offset is far from the resonance of the structure. As Figure 14 shows, the cavity forward and refection signals have extremely similar waveforms, which indicate that the rf power forwarded to the structure is almost fully reflected.
@@ V Conclusion
In this high-power test of the C3 accelerating structure driven by NG-LLRF, the amplitude and phase stability levels have been evaluated comprehensively at different stages of the test setup. The direct loop-back characterization test of the NG-LLRF prototype demonstrated around 70 fs phase jitters and 0.13 % of amplitude fluctuation. This is a critical milestone for the new direct rf sampling technique employed to implement the NG-LLRF platform. The rf pulses are directly generated by the DACs and measured by the ADCs without any analogue rf mixing in higher order Nyquist zones, which can significantly simplify the structure of the LLRF system for new accelerator developments and upgrades for existing ones. For instance, the LCLS-II has 35 cryomodules and each of them requires 4 rf inputs and 2 RF outputs for control and monitoring purposes. The RF control system for LCLS-II requires at least 210 independent RF channels with up and down mixers with local oscillator distributed to each of and discrete data converters with conventional LLRF control platforms. With the NG-LLRF platform, all the analog rf mixers can be eliminated, and the integration level is highly improved. Therefore, the NG-LLRF platform can deliver significant power and footprint reductions, and higher adaptability compared with conventional architectures.
As the up and down conversions of the rf signals are fully realized in digital domain for NG-LLRF, the system can be rapidly configured for different operation modes for rf stations or adapted to accelerators operated at rf frequency lower than 6 GHz by using software-defined methods.The development of the NG-LLRF platform is still in progress. To maximize the adaptability and scalability of the NG-LLRF across applications with different requirements, we are employing modularized approaches to develop the hardware, firmware and software components of the system.
@@
The high-power test with square rf pulse reveals the full filling and dissipation processes and other reflections between the rf components. On the cavity forward signal, there is a decline after the initial ramp up, which is the result of cross coupling between forward and reflection couplers. For accelerators operating with bunch trains, the requirement in the flatness of the pulse top is more stringent. Therefore, the designs and setup of the couplers to sample the rf signals and the connection waveguides should be tuned to minimize the cross coupling. Digital signal processing techniques, such as the subtraction applied in this case, can also be implemented to further improve the measurement accuracy of rf signals at different stages of the structure. For the tests for rf pulse with phase reversal, the high power short pulses have been successfully generated on the reflection signal. Therefore, the NG-LLRF can be employed to drive and control a SLED structure for pulse compression without any analogue modulator or phase shifter. For the tests of rf pulse with a linear phase sweep in 500 ns, the klystron generates the phase ramp with high precision, but rf power forwarded to the cavity has been fully reflected, as the phase ramp is equivalent of drive the structure off the resonance by approximately 2 MHz.